OpenStudy (kanard):

Please help me out. Just the way you got the Problem I dont really need the anwser

OpenStudy (kanard):

@bmk614

563blackghost (563blackghost):

To solve you would basically divide \(\large{2.8}\) and \(\large{1.4}\) BUT you would subtract the exponents of \(\Large{10^{7}}\) and \(\Large{10^{3}}\)... \(\Huge{\frac{2.8}{1.4} \times 10^{7-3}=?}\)

563blackghost (563blackghost):

Were you able to find your answer?